OpenStudy (andriod09):

Guys, need help. last question my my review. totally forgot how to do it, and I can't find it in my notes at all. Help me please: The sum of two number is 27. When 4 times the smaller is subtracted from the larger, the result is 7. Find the numbers.

OpenStudy (andriod09):

Here's the question again: The sum of two number is 27. When 4 times the smaller is subtracted from the larger, the result is 7. Find the numbers.

OpenStudy (anonymous):

\[a+b=27\]\[a-4b=7\]\[a=27-b\], put this value of a in the second equation, you will have: \[27-b-4b=7\]\[27-5b=7\]\[20=5b , b=4\] So, if \[a+b=27 & b = 4, a = 23.\]

OpenStudy (andriod09):

You're just confusing me more. I'm a freshman, and reviewing this for Honours algebra. I've never seen this method at all.

OpenStudy (anonymous):

Ok. You have the two equations from the problem: \[a+b=27\]\[a-4b=7\] A way to solve this is to subtract the first equation of the second: \[a-a=0, b - (-4b) = 5b, 27 - 7 = 20\] You will have: \[5b=20, b=4\] If \[b=4, a=23\]. Where you are having troubles?

OpenStudy (andriod09):

The fact you're using two variables. We don't use that, concidering i'm dyslexic. My teacher teaches me to use \[x_1\space \text{and}\space x_2\]

OpenStudy (anonymous):

\[x _{1}\] and \[x _{2}\] are two variables... Just change \[a=x _{1}\\b=x _{2}\]

OpenStudy (andriod09):

Because to set it up, we would use: (x) ->smaller number (27-x) ->larger number

OpenStudy (anonymous):

Ok, x = smaller number and 27-x = larger number. So, [27 - x] - [4x] = 7 (the large minus four times the smaller). 27 - x - 4x = 7, 27 - 5x = 7 , 20 = 5x , x = 4. So the number are x and 27 - x , 4 and 23.